OpenStudy (anonymous):

Bob and James are finishing the roof of a house. Working alone, Bob can shingle the roof in 10 hours. James can shingle the same roof in 16 hours. How long will it take them working together to shingle the roof? Round your answer to the nearest hundredth if necessary. (1 point) A.13 hours B.26.67 hours C.6.15 hours D.0.04 hours

OpenStudy (anonymous):

Let t = time for both to finish the job 1/10 + 1/16 = 1/t 13/80 = 1/t t = 80/13 ≈ 6.1538.. It would take approximately 6.15 hr working together.

OpenStudy (anonymous):

Thankss (:

OpenStudy (anonymous):

welcome